the intensity of sound wave gets reduced by 20% on passing through a slab .the reduction in intensity on passing through wo consecutive slabs is
 

Grade:9

1 Answers

Biswajit Das
9 years ago
We know that intensity of sound varies, I1= 1.26 I0.
So, 20% of I0will be, (20% of I0) = (20/100) I0= I0/5
Thus the new intensity of sound will be, I1= 1.26 (I0/5) = (1.26/5) I0= 0.252 I0
Therefore from the above observation we conclde that, the new intensity will be0.252 I0.